# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v1

@@ 6. Conclusion
In this paper, we provide an overview of the recent advancements of Artificial Intelligence (AI) models for diagnosing neurodegenerative diseases (NDs) through human gait. Our survey showcases a general process of gait-based NDs diagnosis using AI with focus on five representative NDs: PD, AD, HD, ALS, MS, and systematically reviews all existing studies with an elaboration on their used gait data, AI models, and overall performance. We outline and discuss the current challenges, potential solutions, and promising future directions in this field. Furthermore, we propose a novel research vision on the utilization of emerging 3D skeleton data for human gait representation and AI model development for more efficient NDs diagnosis.

### v2

@@ 6 Conclusion
In this paper, we provide an overview of the recent advancements of Artificial Intelligence (AI) models for diagnosing neurodegenerative diseases (NDs) through human gait. Our survey showcases a general process of gait-based NDs diagnosis using AI with focus on five representative NDs: PD, AD, HD, ALS, MS, and systematically reviews all existing studies with an elaboration on their used gait data, AI models, and overall performance. A novel quality evaluation criterion is also devised to quantitatively assess the quality of surveyed studies. We outline and discuss the current challenges, potential solutions, and promising future directions in this field. Furthermore, we propose a novel research vision on the utilization of emerging 3D skeleton data for human gait representation and AI model development for more efficient NDs diagnosis.